OpenStudy (anonymous):

Hi all L=C+C1+C2+C3+....+Cn C=2*r* Π C1=2*(r-X)* Π C2=2*(r-2X)* Π C3=2*(r-3X)* Π ... Cn=2*(r-nX)*Π How to find n?

OpenStudy (anonymous):

i wish i could help but have no clue haha

OpenStudy (anonymous):

\[(\frac{ C _{n} }{ 2\pi }-r)\frac{ 1 }{ x }=n\] Not sure if this is what your looking for. I just used algebra to make an expression for n.

OpenStudy (anonymous):

equation* not expression.

OpenStudy (amistre64):

i spose L is known?

OpenStudy (amistre64):

\[L=\sum_{i=0}^{n}2\pi(r-ix)\]

OpenStudy (amistre64):

we can pull out constants; and split the other term to start with

OpenStudy (amistre64):

\[L=\sum_{i=0}^{n}2\pi(r-ix)\] \[L=2\pi\sum_{i=0}^{n}(r-ix)\] \[\frac{L}{2\pi}=\sum_{i=0}^{n}(r-ix)\] \[\frac{L}{2\pi}=\sum_{i=0}^{n}r-\sum_{i=0}^{n}ix\] \[\frac{L}{2\pi}=r(n+1)-x\sum_{i=0}^{n}i\] \[\frac{L}{2\pi}=r(n+1)-x\frac{(n+1)(0+n)}{2}\]

OpenStudy (amistre64):

looks like we can multiply thru by 2 and solve the quadritic for n using the quadratic formula if need be

OpenStudy (anonymous):

all but n are known

OpenStudy (amistre64):

then by that last setup, you should be able to figure out the n

OpenStudy (anonymous):



OpenStudy (amistre64):

youre quite welcome :)